OpenStudy (wwe123):

if the ratio of the equation x^2+px+q=0 to the ratio of x^2+lx+m=0,

OpenStudy (wwe123):

Prove that \[ m p^{2}=q l^{2}\]

OpenStudy (wwe123):

sorry

OpenStudy (wwe123):

if the ratio of the equation x^2+px+q=0 is equal to the ratio of equation x^2+lx+m=0,

OpenStudy (anonymous):

U mean ratio of roots?

OpenStudy (wwe123):

yes, but i done the question

OpenStudy (wwe123):

thanks @sauravshakya & @Callisto :)

OpenStudy (wwe123):

Let α and β be the roots of equation x 2 + px + q = 0 and a and b be the roots of equation x 2 + lx + m = 0 Then α + β = – p, αβ = q ......(1) a + b = – l, ab = m ......(2) Given

OpenStudy (wwe123):

@Callisto

OpenStudy (wwe123):

@sauravshakya